OpenStudy (anonymous):

Simplify this expression: 1/tan^2x + cotx tanx I tried to simplify it and came up with 2+tan^2x. Does that seem right to anyone?

OpenStudy (anonymous):

no... it should be cos^2 x

OpenStudy (anonymous):

cosec^2x is the answer

OpenStudy (anonymous):

+cotx tanx is wid tan^2x ?? like in denominator?

OpenStudy (anonymous):

Interesting...how did you both arrive there? I'll try at it to see....

OpenStudy (anonymous):

see cotx tanx= tanx/tanx=1 now.. 1/(tan^2x+1)=1/(sec^2x)= cos^2x

OpenStudy (anonymous):

No, +cotx tanx is seperate.

OpenStudy (anonymous):

so its cot^2x+1 = cosec^2x

OpenStudy (anonymous):

1/tan^2x +1 1+tan^2x/tan^2x sec^2x/tan^2x cosec^2x

OpenStudy (anonymous):

So then it simplies to cosxsec^2x?

OpenStudy (anonymous):

OK now I'm getting that it equals 1......1/tan^2x + cotxtanx. Then 1/tan^2x + 1. Then, 1/tan^2x +tan^2x/tan^2x, equals 1+tan^2x/tan^2x, which simplifies to 1.

OpenStudy (rogue):

Is this your problem?\[\frac {1}{\tan^2x} + \cot x \tan x\]If so, that's csc^2 x.

OpenStudy (rogue):

\[\frac {1}{\tan^2x} + \cot x \tan x = \cot^2 x + 1 = \csc^2 x\]

OpenStudy (anonymous):

Yeah, that's my problem @Rogue ! I see now, that makes sense!

OpenStudy (rogue):

Glad I could help =)

OpenStudy (anonymous):

:)